2. Suppose you are asked to design a space station that is 1.0 km in diameter, and is spun fast enough so that a resident standing within it at the outer circumference of the station would experience earth normal gravity. How fast (in rotations per minute, rpm) would you have to spin the space station?
1
Expert's answer
2020-03-23T10:47:20-0400

The centripetal acceleration on the resident standing within it at the outer circumference of the station


"a=\omega^2R=g."

Hence


"\omega=\sqrt{\frac{g}{R}}=\sqrt{\frac{9.8}{500}}=0.14\:\rm rad/s=1.34\:\rm rpm."